a lump of steel of mass 8 kg at 1000 k is dropped in 80 kg of oil at 300 k. calculate the entropy change of steel the oil and the universe. take specific heats of steel and oil as 0.5 kj/kg K and 3.5 Kj /kg K respectively


1
Expert's answer
2021-11-24T12:00:41-0500

"m_s = 8 \;kg \\

m_o = 80 \; kg \\

T_1 = 1000 \;K \\

T_2 = 300 \;K \\

c_{ps} = 0.5 \; kJ/kg \cdot K \\

c_{po} = 3.5 \; kJ /kg \cdot K"

Heat lost (steel) = Heat gain (Oil)

"m_s c_{ps}(T_1 -T) = m_o c_{po} (T- T_2) \\

8 \times 0.5 \times (1000 -T) = 80 \times 3.5 (T- 300) \\

4000 -4T = 280T -84000 \\

88000 = 284T \\

T = 309.8 \;K"

Entropy change

System: steel "= m_s c_{ps} ln(\frac{T}{T_1})"

"= 8 \times 0.5 \times ln(\frac{309.8}{1000}) = -4.68 \; kJ/K"

Surrounding: Oil "= m_o c_{po} ln(\frac{T}{T_2})"

"= 80 \times 3.5 \times ln(\frac{309.8}{300}) = 9.00 \; kJ/K"

Universe = System – Surrounding

"= -4.68 + 9.00 = 4.32 \;kJ/K"